OpenStudy (anonymous):

I need to find cos(2theta) if sin(theta)=1/5

OpenStudy (anonymous):

Find theta by taking inverse sin of .2...then compute cos(2*theta)

OpenStudy (anonymous):

how do I go about getting the inverse?

OpenStudy (anonymous):

Calculator would be best bet...or maybe a lookup table. I'm sure google would do it for you

OpenStudy (anonymous):

Yeah, I don't think they want them using calculators for this problem...

OpenStudy (anonymous):

yeah it's a take home quiz and we can get help but we aren't supposed to use calculators

OpenStudy (anonymous):

try this : \[\sin(\theta) = 1/5\] so opposite =1 hypotenuse =5

OpenStudy (anonymous):

adjacent = ?

OpenStudy (anonymous):

\[\sqrt{25-1}\]

OpenStudy (anonymous):

Ah yes good thinking

OpenStudy (anonymous):

now look at your trig identities for \[\cos 2 \theta\]

OpenStudy (anonymous):

one of them is going to be easy to use here... since you know \[\cos \theta = (\sqrt{25-1} )/ 5\]

OpenStudy (anonymous):

we haven't gone over identities yet. . .

OpenStudy (anonymous):

you sure?

OpenStudy (anonymous):

pretty sure, I am looking through my notes now. This is a retake quiz cause nobody aced it the first time and it's only three questions so I think the teacher may have forgotten to tell us something vital.

OpenStudy (anonymous):

it's not just \[2\times(\sqrt{25-1})\] is it?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

didn't think so. . .Could you teach me the identity?

OpenStudy (anonymous):

trying to find another way of doing this geometrically, if you really aren't supposed to be using identities...

OpenStudy (anonymous):

do you have a unit circle with (1/5, 1) as one of the angles? That's the easiest way...

OpenStudy (anonymous):

hey I found some identity stuff on google, Do I have to use \[1-2\sin^2(x)\]? and everything else that goes with it?

OpenStudy (anonymous):

"one of the points for which an angle is given" is what I should've said.

OpenStudy (anonymous):

Yeah, man you can use that, but if it's not what your teacher wants you to use... you might not get credit.

OpenStudy (anonymous):

ok I will go for it. so do I just plug 1/5 in for x?

OpenStudy (anonymous):

nope plug 1/5 in for sin(theta)

OpenStudy (anonymous):

\[1-2(1/5)^2\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Yeah, I was thinking of using an identity for cos(2*theta) that had cosine terms in it... but your way is easier... but are you sure you're allowed to use identities to solve this problem?

OpenStudy (anonymous):

yeah I'm pretty sure, it says we will have to use a formula first as a hint on the problem

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So from that formula how do I get to \[\cos(2\theta)\]?

OpenStudy (anonymous):

you already did \[\cos 2\theta =1 -2\sin ^{2} \theta\]

OpenStudy (anonymous):

so that's it? I can just have that as my answer?

OpenStudy (anonymous):

calculate \[1-2(1/5)^{2}\]

OpenStudy (anonymous):

what is it?

OpenStudy (anonymous):

.92 or 23/25

OpenStudy (anonymous):

that's it

OpenStudy (anonymous):

My teacher said we don't have to fully simplify the answers and its a non calculator quiz. So that being said, is the unsimplified version the best answer to the problem?

OpenStudy (anonymous):

23/25 then

OpenStudy (anonymous):

ok sweet thanks for the help! Can I get your help on one more problem? or at can you least make sure I'm doing it right?